# Novel Lower Bounds on $M/G/k$ Scheduling

### v2

@@ 11. Empirical and Numerical Results
In this section, we illustrate the effectiveness of our method by presenting empirical and numerical results on our novel lower bounds on the mean response time in the $M/G/k$ system under an arbitrary scheduling policy, based on the ISQ-$k$ system.
11.1. Total work
Recall that we exactly characterize total mean work in the ISQ-2 system in Proposition 7.3. We prove in Theorem 4.3 that mean work in the ISQ-2 lower bounds mean work in the $M/G/2$ system under an arbitrary scheduling policy. Here, we numerically and empirically verify this result for several scheduling policies.
Figure 3 plots the ratios between the empirical mean total work of an $M/G/2$ system under several scheduling policies, as compared to the exact mean work in the matching ISQ-2 system. We numerically and empirically evaluate these policies under an exponential job size distribution. For FCFS-2 scheduling, the exact mean total work of an $M/G/2$ with exponential job size distribution is known. For SRPT-2 and Longest Remaining Processing Time (LRPT-2) scheduling, we evaluate the policies empirically.
Figure 3 shows that the total mean work of the ISQ-2 system serves as a lower bound for the total work of the $M/G/2$ system under the FCFS, SRPT, and LRPT policies, confirming our result. In our exploration, it appears that the LRPT policy minimizes the total work of an $M/G/2$ system, across all scheduling policies, making this lower bound especially important.
We also compare the ISQ-2 lower bound to the prior resource-pooled $M/G/1$ lower bound, note that in an $M/G/1$, scheduling policy does not affect mean work. We see that the total work of ISQ-2 is significantly larger than the total work of the resource-pooled $M/G/1$ system, particularly for low to medium load, indicating that it is a much stronger bound.
Figure 3. Ratios of the expected total work of $M/G/2$ under different scheduling policies compared to ISQ-2 and resource-pooled $M/G/1$.
11.2. Response time bounds
We now illustrate the quality of the lower bounds on mean response time achievable by adding our lower bounds on mean response time under arbitrary scheduling policies, namely Theorem 9.4 and 9.6 into our WINE framework (see Section 10). In Figure 4, we plot the ratios of the WINE-2, WINE-3, and WINE-4 bounds defined in Section 10 compared to the previous naive lower bounds, namely the resource-pooled SRPT-1 and the mean service time bounds. We consider two servers and an exponential job size distribution.
We observe that the WINE-2 bound, which combines naive lower bounds with our WINE framework significantly improves upon the existing naive response time lower bounds, achieving over a 20% improvement for certain loads. By incorporating the ISQ-$k$ lower bounds, the WINE-3 and WINE-4 bounds offer an even larger improvement. They begin to outperform the baseline bound at much smaller loads, around $\rho=0.3$, and reach improvements of over 25%.
Figure 4. Ratios of WINE bounds over naive lower bounds for $M/G/2$ under arbitrary scheduling. “Baseline” is the maximum of $E[T^{SRPT\text{-}1}]$ and $2E[S]$.
To more precisely measure the degree of improvement of our ISQ-based lower bounds beyond what we already achieved with WINE-2, in Figure 5 we plot the improvements of WINE-3 and WINE-4 over WINE-2 for two different job size distributions with differing variability. In Figure 5(a), we have a low-variability distribution, an Erlang-3 job size distribution with a squared coefficient of variation $C^{2}=\mathrm{Var}(S)/\mathbb{E}[S]^{2}=1/3$. In Figure 5(b), we have a high-variability distribution, a 2-branch hyperexponential job size distribution, with $C^{2}=5$.
(a) Erlang-3, $C^{2}=1/3$.
(b) Hyperexponential, $C^{2}=5$.
Figure 5. Ratios of improvement of the WINE 3 and WINE 4 lower bounds over the baseline WINE-2 lower bound for various job size distributions and $k=2$ servers.
@@ 11.2. Response time bounds
Two important observations become apparent: First, for the lower variability distribution in Figure 5(a), the greater the improvement of the ISQ-$k$ based lower bounds over WINE-2, reaching over $8\%$ for the Erlang-3 distribution. One intuitive explanation for this phenomenon is that the ISQ-$k$ system takes an optimistic view of work completion, focusing on the case where all jobs finish at the same time, so no servers are wasted prematurely. This optimistic view is most accurate when jobs have similar sizes, as in a low variability setting. Secondly, for the high variability job size distributions Figure 5(b), the AR-ISQ-2 based WINE-4 bound which incorporates recycling information offers a much greater improvement over WINE-2 than WINE-3 alone, justifying the need for Theorem 7.6. This makes sense, as the larger, recycled jobs make up more of the load at intermediate thresholds under job size distributions with high variability.
@@ 12. Conclusion
We present the first nontrivial lower bounds on the mean response time of the $M/G/k$ system under arbitrary scheduling policies. We introduce the novel ISQ-$k$ queue which bridges the multiserver and single-server queues and helps us lower bound the mean relevant work of an $M/G/k$ system under arbitrary scheduling policies.
We introduce the novel DiffeDrift extension to the drift method to characterize the work of the increasing speed queue, introducing a new technique where we derive test functions via solutions of differential equations. Empirically, our lower bounds improve upon previous naive lower bounds for a wide range of loads, with the most significant improvement observed under moderate load conditions.
One direction for future work lies in proving 9.7 on jumps due to recycling for $k>2$ servers. More ambitiously, one could try to derive similar lower bounds on the mean response time for the $G/G/k$ queue. While this paper relies on the Poisson arrival process, the underlying ISQ-$k$ system and the DiffeDrift method have the potential to be more broadly applicable. A final potential direction is to prove even tighter bounds by first proving that the LRPT scheduling policy minimizes remaining total work in the $M/G/k$.
@@ Appendix A Proofs for Section 5
00 0 $=\dfrac{1}{t}\int_{\mathcal{S}}\left(\int_{\mathcal{S}}g(y)P(t,x,dy)-g(x)\right)\pi(dx)=\lim_{t\to 0}\int_{\mathcal{S}}\frac{1}{t}\left(\int_{\mathcal{S}}g(y)P(t,x,dy)-g(x)\right)\pi(dx).$
@@
$M/G/1$ queues with rest periods are studied in [28]. Equation 15 of [28] characterizes the expected second moment of waiting time of an $M/G/1$ queue with rest periods. In particular, this expectation is finite if $E[S^{3}]<\infty$ and $\mathbb{E}[T_{0}^{3}]<\infty$. Note that the sum of $k$ exponential distributions has finite moments of all orders, so the lemma holds by our assumption that $\mathbb{E}[S^{3}]<\infty$. Note that [28] assumes that $T_{0}$ is independent of the arrival process, which is not the case in our setting. However the proof of Equation 15 proceeds by first characterizing the number of arrivals during the rest period, before proceeding with the rest of the proof. In our system, that number of arrivals always $k$, simplifying Equation 7 of [28], and remainder of the proof holds unchanged. ∎
See 5.4
@@
To illustrate the high-level idea, we first consider the case where the speed $i$ is 1. By the definition of the Poisson arrival process, for $t<w$,
@@
Before we continue, note that to ensure uniform convergence, we must also consider the case that $t>w$, for $w$ close to 0. In this case, the only significant change is the at the $N_{t}=0$ term results in a state after time $t$ of $(0,0)$, rather than $(w-t,1)$. By assumption, $\lim_{w\to 0^{+}}g(w,1)=g(0,0)$, so the proof proceeds unchanged.
@@
By the independence of $S$ and $N_{t}$, it is easy to check that $\mathbb{E}[g(w-t+N_{t}\cdot S,1)\,|\,N_{t}>1]<\infty$ given the assumption that $g(w,i)=w^{2}+c(w,i)$ where $|c(w,i)|\leq C_{1}w+C_{2}$. Therefore, we have, $\lim_{t\to 0}\mathbb{E}[B(t,W,1)]=0.$ We now want to show that $A(t,\cdot,1)$ converges uniformly to some limit as $t\to 0$, namely, $G\circ g$. We first consider the term $-g(w,1)/t+g(w-t,1)/t$. Since $g$ is twice-differentiable for all $w\in\mathbb{R}_{+}$, by Taylor expansion,
$g(w,1)-g(w-t,1)=g^{\prime}(w,1)t-R_{1}(w,t)=g^{\prime}(w,1)t-g^{\prime\prime}(c,1)\frac{t^{2}}{2}$ (A.3)
Recall that by our assumption, $c(w,i)$ has a bounded second derivative and $g(w,i)$ also has a bounded second derivative. Therefore, $\dfrac{1}{t}(-g(w,1)+g(w-t,1))=-g^{\prime}(w,1)+\frac{g^{\prime\prime}(c,1)}{2}t\to-g^{\prime}(w,1)$ uniformly for all $w$. Next, we want to show that $\lambda\mathbb{E}[g(S+w-t)]-\lambda g(w-t)\to\lambda\mathbb{E}[g(S+w)]-\lambda g(w)$ uniformly. Let $\hat{g}(w-t,1)$ be defined as this difference, namely $\mathbb{E}[g(S+w-t,1)]-g(w-t,1)$. Then again since $\hat{g}$ is smooth, we have by Taylor’s theorem
@@
When the speed of the system is less than 1, the only necessary changes lie in Equations A.3 and A. Eq. A.3 becomes $g(w,i/k)-g(w-it/k,i/k)=g^{\prime}(w,i/k)it/k-\bar{R}_{1}(w,t),$ for some remainder term $\bar{R}_{1}$. We now invoke our bounded second derivative assumption to bound $\bar{R}_{1}(w,t)$ in the same manner as $R_{1}(w,t)$ above. Dividing by $t$ and taking the limit as $t\to 0$ yields the desired limit $g^{\prime}(w,i/k)\frac{i}{k}$. Similarly, Appendix A can now be written as,
@@ Appendix B Proofs for Section 9
$g_{k}(\cdot,\cdot)$ defined by Equation 9.1 satisfies the assumptions of Proposition 5.2 and Lemma 5.4. In particular, $\lim_{w^{+}\to 0}g_{k}(w,i)=0$. Note also that $g_{k}(\cdot,\cdot)\geq 0$.
@@
Moreover, one can see that from the form of Equation 9.1 that each $u_{q}$ is a sum of linear and negative exponential terms in $w$. In particular, in the integrand, $\mathbb{E}[u_{q+1}(S+y)]$ preserves this structure during each recursive step, see, e.g., the ISQ-3 test functions (C.1) and (C.2) given in Appendix C. Therefore, $u_{q}$ must satisfy the assumptions of Proposition 5.2 and Lemma 5.4.
@@
For any arbitrary job size distribution $S$ such that $E[S^{3}]$ is finite, the drift at speed 0 is $G\circ g_{k}(0,0)=\lambda\mathbb{E}[S+u_{1}(S)]$ and the drift at all other speeds $i\geq 1/k$ is given by $G\circ g_{k}(w,i)=\lambda\mathbb{E}[S]-1$ for all $i\geq 1/k$.
@@
Clearly, the drift at speed $1$ is given by $G\circ g_{k}(w,1)=\lambda\mathbb{E}[S]-1$ and the drift at speed 00 is given by $G\circ g_{k}(0,0)=\lambda\mathbb{E}[S+u_{1}(S)]$. For all other speed $q/k$, the drift is given by,
@@
For any arbitrary job size distribution $S$ such that $E[S^{3}]$ is finite, the drift at speed 00 is given by $G\circ h_{k}(0,0)=\lambda\mathbb{E}[S^{2}+v_{1}(S)]$ and the drift at all other speed $\geq 1/k$ is given by $G\circ h_{k}(w,i)=\lambda\mathbb{E}[S^{2}]+2w(\lambda\mathbb{E}[S]-1)$.
@@
Clearly, the drift at speed $1$ is given by $G\circ h_{k}(w,1)=\lambda\mathbb{E}[S^{2}]+2w(\lambda\mathbb{E}[S]-1)$ and the drift at speed 00 is given by $G\circ h_{k}(0,0)=\lambda\mathbb{E}[S^{2}+v_{1}(S)]$. For all other speed $q/k$, the drift is given by,
@@
For any arbitrary truncated job size distribution $S_{x}$ and arrival rate $\lambda_{x}$, the drift is given by
@@
Recall that $C_{k}(x,\lambda_{x})$ is defined to be $C_{k}(x,\lambda_{x})=\frac{2}{k}\cdot\frac{\lambda_{x}\mathbb{E}[v_{1}(S_{x})]}{2+2\lambda_{x}\mathbb{E}[u_{1}(S_{x})]}$. Recall also that by Proposition 9.3, we have the following alternative expression for mean relevant work for the same arrival rate $\lambda_{x}$ and truncated job size distribution $S_{x}$:
@@
Therefore, we can write the drift at speed 0 as $G\circ h_{k,x}(0,0)=\lambda_{x}\mathbb{E}[S_{x}^{2}]+kC_{k}(x,\lambda_{x})$. We are now ready to state the general version of the AR-ISQ-$k$ lower bound. This is the second lower bound of Theorem 4.2.
See 9.6
@@
The proof of this theorem follows the same argument as the proof of Theorem 7.6. Let $\mathbb{E}_{r}[\cdot]$ denote the Palm expectation taken over the moments when the arbitrary arrival of jobs of size $x$ occur.
Let $J_{x}(w,i):=h_{k,x}(w+x,\min(i+1/k,1))-h_{k,x}(w,i)$ denote the increase in the test function due to the arrival of a size-$x$ job. Then $\mathbb{E}_{r}[J_{x}(W_{x},I)]$ denotes the mean size of the recycling jump in stationarity. Thus, by Lemma 5.6,$\mathbb{E}[G\circ h_{k,x}(W_{x},I)]=\mathbb{E}[\text{stochastic drift}]+(\lambda-\lambda_{x})\mathbb{E}_{r}[J_{x}(W_{x},I)]$. By Lemma B.7 we can write the expected stochastic drift as,
$\mathbb{E}[\text{stochastic drift}]$ $=\lambda_{x}\mathbb{E}[S_{x}^{2}]+2\mathbb{E}[W_{x}](-1+\rho_{x})+kC_{k}(x,\lambda)\sum_{i=0}^{k}\frac{k-i}{k}\mathbb{P}(I^{r}=i/k)$
@@
$\mathbb{E}[W_{x}^{arb{\text{-}}ISQ}]$ $=\frac{\lambda_{x}E[S_{x}^{2}]}{2(1-\rho_{x})}+\dfrac{kC_{k}(x,\lambda_{x})(1-\rho_{\mkern 1.5mu\overline{\mkern-1.5mux\mkern-1.5mu}\mkern 1.5mu})}{2(1-\rho_{x})}+\frac{(\lambda_{x}-\lambda_{x})\mathbb{E}_{r}[J_{x}(W_{x},I)]}{2(1-\rho_{x})}$
$=\dfrac{\lambda_{x}\mathbb{E}[S_{x}^{2}]}{2(1-\lambda_{x}\mathbb{E}[S_{x}])}+\dfrac{\lambda_{x}\mathbb{E}[v_{1}(S_{x})]}{2+2\lambda_{x}\mathbb{E}[u_{1}(S_{x})]}\cdot\dfrac{1-\rho_{\bar{x}}}{1-\rho_{x}}+\frac{(\lambda-\lambda_{x})\mathbb{E}_{r}[J_{x}(W_{x},I)]}{2(1-\rho_{x})},$
where we use the definition of $C_{k}(x,\lambda)$ in the second equality. Recall the definition of an increasing speed queue. The current speed $i$ bounds the number of jobs in the system, which in turn bounds the total work, which consists of at most $ki$ jobs, each with size at most $x$. Therefore, for $10<i<$ 1, the system can only visit states $(w,i)$ where $w\leq kix$.

### v3

@@ 12. Empirical and Numerical Results
In this section, we demonstrate the effectiveness of our method by presenting empirical and numerical results for our novel lower bounds on the mean response time in the $M/G/k$ system under arbitrary scheduling policies. We focus on our performance metric, the Uncertainty Improvement Ratio (UIR) which we define in Section 12.1, and examine its behavior under a variety of settings.
We investigate the impact of changing the number of servers in Section 12.2, of job size variability in Section 12.3, and finally compare the UIR of our lower bounds against recently proposed upper bounds in the literature in Section 12.4.
12.1. Uncertainty Improvement Ratio
To quantify the quality of our lower bounds on mean response time (Theorems 4.1 and 4.2) we use the Uncertainty Improvement Ratio (UIR).
This metric captures the shrinkage of the uncertainty region where the optimal policy may lie. It measures how much of this region is closed by a new lower bound relative to the best prior lower bound.
Definition 12.1.
The Uncertainty Improvement Ration of a response time lower bound $\mathbb{E}[T_{\text{novel-lower}}]$ relative to a pair of prior response time lower and upper bounds is given by
$\mathrm{UIR}=\frac{\mathbb{E}[T_{\text{novel-lower}}]-\mathbb{E}[T_{\text{lower}}]}{\mathbb{E}[T_{\text{upper}}]-\mathbb{E}[T_{\text{lower}}]}.$
Here $\mathbb{E}[T_{\text{lower}}]$ is a prior lower bound and $\mathbb{E}[T_{\text{upper}}]$ is a prior upper bound on the response time of $M/G/k$. We will use the simulated SRPT-$k$ response time as the upper bound. By definition, $\mathrm{UIR}\leq 100\%$, with $\mathrm{UIR}=0\%$ indicating no improvement over the prior bound and $\mathrm{UIR}=100\%$ indicating full closure of the uncertainty region.
12.2. Varying number of servers $k$
In this section, we examine the UIR of the ISQ-$k$ system. We study how the UIR of the MixEx bound relative to the naive lower bounds changes as $k$ increases. We also examine how the UIR of our novel bounds relative to the MixEx bound varies with $k$. Computing the response time requires numerically solving Equations 4.9 and 4.10. The explicit expressions for the functions $v_{1}$ and $u_{1}$ appearing in those equations are given by Equations 8.1 and 8.2 for $k=2$ and also in Appendix C for $k=3,4,$ and $5$. This numerical integration, performed using Mathematica or SciPy, does not present computational difficulties. We discuss how these functions could be computed for larger $k>5$ using a specialized computational program in Section 11.
Figure 3. UIR of MixEx bound relative to naive bounds in the $M/G/k$ setting with exponential job sizes. Load ranges from 0.3 to 0.95 with five million simulated arrivals per load.
The UIR of the MixEx bound relative to the naive lower bounds is shown in Figure 3. For $k=2,3,4$ and $5$, we observe that the maximum UIR of the MixEx bound increases as $k$ increases. However, the improvement only begins to deviate from zero at progressively higher loads levels as $k$ increases, with the transition occurring approximately at $\rho\approx 1-1/k$.
Figure 4. UIR of novel bounds relative to MixEx bound in the $M/G/k$ setting with exponential job sizes. Load ranges from 0.3 to 0.95 with five million simulated arrivals per load.
The UIR of our novel lower bounds relative to the MixEx bound is shown in Figure 4.
These bounds begin closing the uncertainty gap at significantly lower loads than MixEx and sustain meaningful improvement even as $\rho$ approaches 0.95. However, in contrast to the MixEx bound, the maximum UIR of the ISQ and ISQ-Recycling bounds decreases as $k$ increases. Note that, Figure 4 should be viewed as an improvement upon the improvements already shown in Figure 3. Taken together, the two bounds result in a maximum UIR improvement of roughly 60% for all $k=2$ to 5, resulting in a substantial reduction in the uncertainty region at the optimal policy. This improvement is demonstrated across a wide range of load.
While the ISQ and ISQ-Recycling bounds are comparable for lower loads in this setting, the ISQ-Recycling bound provides a significant advantage at higher loads. As we show in the next subsection, this advantage becomes even more pronounced when the job size distribution has large variability (see Figure 6).
12.3. Job size variability
In this section, we study the improvements of the ISQ and ISQ-Recycling bounds over the MixEx bound for two families of job size distributions with differing variability, in the setting of two servers.
In Figure 5, we consider a low variability case with three uniform job size distributions, where the squared coefficient of variation is $C^{2}=\mathrm{Var}(S)/\mathbb{E}[S]^{2}\in\{0.01,0.05,0.2\}$. In Figure 6, we examine a high variability case with three two-branch hyperexponential job size distributions, where $C^{2}\in\{2,3,5\}$.
Figure 5. UIR relative to the MixEx bound for the ISQ and ISQ-Recycling bounds in the $M/G/2$ setting with Uniform job size distributions of various $C^{2}$. Load ranges from 0.1 to 0.95 with five million simulated arrivals per load.
Important observations become apparent in both the high and low variability settings. First, for the low variability distribution, the ISQ-$k$ based lower bounds achieve substantially greater improvement over MixEx, reaching a UIR of 62% for the uniform distribution with $C^{2}=0.05$. Moreover, the improvement spreads broadly across the load range. Finally, the UIR continues to rise as the job size distribution approaches a deterministic distribution, approaching a limit just above 70%.
An intuitive explanation is that, in low variability settings, busy periods are short and have similar lengths. Without large jobs to trigger long busy periods, the ISQ-$k$ system spends most of its time operating below a speed of 1. This makes ISQ-$k$ behave more similarly to a true $M/G/k$ system, allowing the ISQ and ISQ-Recycling bounds to yield substantially larger improvements over MixEx.
Figure 6. UIR relative to the MixEx bound for the ISQ and ISQ-Recycling bounds in the $M/G/2$ setting with hyperexponential job size distributions of various $C^{2}$. Load ranges from 0.4 to 0.95 with five million simulated arrivals per load.
@@ 12.3. Job size variability
Secondly, for the high variability job size distributions, the Rec-ISQ-2 based ISQ-Recycling bound which incorporates recycling information offers a much greater improvement over MixEx than ISQ bound alone, justifying the need for Theorem 8.6. This makes sense, as the larger, recycled jobs make up more of the relevant work at intermediate thresholds under job size distributions with high variability. We also observe that the maximum UIR decreases as variability increases and the peak shifts to progressively higher load levels.
12.4. Upper bound comparison
In this subsection, we compare the UIR of our novel lower bounds relative to the MixEx bound against the UIR of a recently introduced scheduling policy.
This upper bound corresponds to the the recently introduced SRPT-Except-$k+1$ (SEK) policy (Grosof2024BoMS and Grosof2026), which is proven to outperform SRPT-$k$ at all load levels and aims to tighten upper bounds on the mean response time of the optimal $M/G/k$ scheduling policy.
In particular, we empirically study the UIR of the Practical SEK policy, a variant of the core SEK policy which has been empirically shown to improve on SRPT-$k$.
We consider the setting of two servers with an exponential job size distribution. As shown in Figure 7, uur best lower bound, the ISQ-Recycling bound, achieves a maximum UIR of approximately 34%, whereas the Practical-SEK policy attains a UIR of only about 3%. See Appendix D for additional details on the Practical-SEK policy and the computational experiments.
Thus, our lower bounds close the uncertainty region far more dramatically from below than Practical-SEK does from above.
Figure 7. UIR relative to the SRPT-$k$ and the MixEx bound for the ISQ-Recycling and Practical-SEK bounds in the $M/G/2$ setting with $\text{Exp}(1)$ job sizes. Load ranges from 0.3 to 0.95 with five million simulated arrivals per load.
@@ 13. Conclusion
We present the first nontrivial lower bounds on the mean response time of the $M/G/k$ system under arbitrary scheduling policies.
We start by applying the WINE combination to existing lower bounds, producing the MixEx bound, which substantially tightens the uncertainty region of the optimal policy, though it remains loose at intermediate loads.
Our main contributions are the ISQ-$k$ system and the DiffeDrift method, which together yield two novel lower bounds on response time, the ISQ and ISQ-Recycling bounds. Our ISQ and ISQ-Recycling bounds achieve significantly tighter results than MixEx bound.
Empirically, our lower bounds improve upon the prior naive bounds and the MixEx bound across a wide range of loads, with the most significant improvement appearing under moderate load conditions. For exponential job size distributions, the ISQ-Recycling bound achieves a maximum Uncertainty Improvement Ration (UIR) of 34% on top of the MixEx bound, which itself provides a substantial improvement over the naive bounds. Moreover, for high variability job sizes, the ISQ-Recycling bound exhibits substantially higher UIR than the ISQ bound.
Future directions include proving 10.7 which concerns the behavior of jumps due to recycling for $k>2$ servers, extending our framework to the setting of unknown job sizes as discussed in Section 11, generalizing to the $G/G/k$ model, and analyzing the asymptotic scaling of our bounds as $k$ grows large.
@@ Appendix A Proofs for Section 7
$0$ $=\dfrac{1}{t}\int_{\mathcal{S}}\left(\int_{\mathcal{S}}g(y)P(t,x,dy)-g(x)\right)\pi(dx)=\lim_{t\to 0}\int_{\mathcal{S}}\frac{1}{t}\left(\int_{\mathcal{S}}g(y)P(t,x,dy)-g(x)\right)\pi(dx).$
@@
$M/G/1$ queues with rest periods are studied in Scholl1983O. Scholl1983O characterize the expected second moment of waiting time of an $M/G/1$ queue with rest periods. In particular, this expectation is finite if $E[S^{3}]<\infty$ and $\mathbb{E}[T_{0}^{3}]<\infty$. Note that the sum of $k$ exponential distributions has finite moments of all orders, so the lemma holds by our assumption that $\mathbb{E}[S^{3}]<\infty$. Note that Scholl1983O assume that $T_{0}$ is independent of the arrival process, which is not the case in our setting. However the proof of [Scholl1983O, Equation (15)] proceeds by first characterizing the number of arrivals during the rest period, before proceeding with the rest of the proof. In our system, that number of arrivals always $k$, simplifying Scholl1983O, and remainder of the proof holds unchanged. ∎
See 7.4
@@
To illustrate the high level idea, we first consider the case where the speed $i$ is 1. By the definition of the Poisson arrival process, for $t<w$,
@@
Before we continue, note that to ensure uniform convergence, we must also consider the case that $t>w$, for $w$ close to 0. In this case, the only significant change is the at the $N_{t}=0$ term results in a state after time $t$ of $(0,0)$, rather than (wt,1)(w-t,1). By assumption, $\lim_{w\to 0^{+}}g(w,1)=g(0,0)$, so the proof proceeds unchanged.
@@
By the independence of $S$ and $N_{t}$, it is easy to check that $\mathbb{E}[g(w-t+N_{t}\cdot S,1)\,|\,N_{t}>1]<\infty$ given the assumption that $g(w,i)=w^{2}+c(w,i)$ where $|c(w,i)|\leq C_{1}w+C_{2}$. Therefore, we have, $\lim_{t\to 0}\mathbb{E}[B(t,W,1)]=0.$ We now want to show that $A(t,\cdot,1)$ converges uniformly to some limit as $t\to 0$, namely, $G\circ g$. We first consider the term g(w,1)/t+g(wt,1)/t-g(w,1)/t+g(w-t,1)/t. Since $g$ is twice-differentiable for all $w\in\mathbb{R}_{+}$, by Taylor expansion,
$g(w,1)-g(w-t,1)=g^{\prime}(w,1)t-R_{1}(w,t)=g^{\prime}(w,1)t-g^{\prime\prime}(c,1)\frac{t^{2}}{2},$ (A.3)
where $R_{1}$ is some suitable remainder term. Recall that by our assumption, $c(w,i)$ has a bounded second derivative and $g(w,i)$ also has a bounded second derivative. Therefore, $\dfrac{1}{t}(-g(w,1)+g(w-t,1))=-g^{\prime}(w,1)+\frac{g^{\prime\prime}(c,1)}{2}t\to-g^{\prime}(w,1)$ uniformly for all $w$. Next, we want to show that $\lambda\mathbb{E}[g(S+w-t)]-\lambda g(w-t)\to\lambda\mathbb{E}[g(S+w)]-\lambda g(w)$ uniformly. Let $\hat{g}(w-t,1)$ be defined as this difference, namely $\mathbb{E}[g(S+w-t,1)]-g(w-t,1)$. Then again since $\hat{g}$ is smooth, we have by Taylor’s theorem
@@
When the speed of the system is less than 1, the only necessary changes lie in Equations A.3 and A. Equation A.3 becomes $g(w,i/k)-g(w-it/k,i/k)=g^{\prime}(w,i/k)it/k-\bar{R}_{1}(w,t),$ for some remainder term $\bar{R}_{1}$. We now invoke our bounded second derivative assumption to bound $\bar{R}_{1}(w,t)$ in the same manner as $R_{1}(w,t)$ above. Dividing by $t$ and taking the limit as $t\to 0$ yields the desired limit $g^{\prime}(w,i/k)\frac{i}{k}$. Similarly, Appendix A can now be written as,
@@ Appendix B Proofs for Section 9
$g_{k}(\cdot,\cdot)$ defined by Equation 10.1 satisfies the assumptions of Proposition 7.2 and Lemma 7.4. In particular, $\lim_{w\to 0^{+}}g_{k}(w,i)=0$. Note also that $g_{k}(\cdot,\cdot)\geq 0$.
@@
Moreover, one can see that from the form of Equation 10.1 that each $u_{q}$ is a sum of linear and negative exponential terms in $w$. In particular, in the integrand, $\mathbb{E}[u_{q+1}(S+y)]$ preserves this structure during each recursive step, see, e.g., the ISQ-3 test functions given in Appendix C. Therefore, $u_{q}$ must satisfy the assumptions of Proposition 7.2 and Lemma 7.4.
@@
For any arbitrary job size $S$ such that $E[S^{3}]$ is finite, the drift at speed 0 is $G\circ g_{k}(0,0)=\lambda\mathbb{E}[S+u_{1}(S)]$ and the drift at all other speeds $i\geq 1/k$ is given by $G\circ g_{k}(w,i)=\lambda\mathbb{E}[S]-1$ for all $i\geq 1/k$.
@@
Clearly, the drift at speed $1$ is given by $G\circ g_{k}(w,1)=\lambda\mathbb{E}[S]-1$ and the drift at speed 0 is given by $G\circ g_{k}(0,0)=\lambda\mathbb{E}[S+u_{1}(S)]$. For all other speed $q/k$, the drift is given by,
@@
For any arbitrary job size $S$ such that $E[S^{3}]$ is finite, the drift at speed 0 is given by $G\circ h_{k}(0,0)=\lambda\mathbb{E}[S^{2}+v_{1}(S)]$ and the drift at all other speed $\geq 1/k$ is given by $G\circ h_{k}(w,i)=\lambda\mathbb{E}[S^{2}]+2w(\lambda\mathbb{E}[S]-1)$.
@@
Clearly, the drift at speed $1$ is given by $G\circ h_{k}(w,1)=\lambda\mathbb{E}[S^{2}]+2w(\lambda\mathbb{E}[S]-1)$ and the drift at speed 0 is given by $G\circ h_{k}(0,0)=\lambda\mathbb{E}[S^{2}+v_{1}(S)]$. For all other speed $q/k$, the drift is given by,
@@
For any arbitrary truncated job size $S_{x}$ and arrival rate $\lambda_{x}$, the drift at $h_{k,x}$ is given by
@@
Recall that $C_{k}(x,\lambda_{x})$ is defined to be $C_{k}(x,\lambda_{x})=\frac{2}{k}\cdot\frac{\lambda_{x}\mathbb{E}[v_{1}(S_{x})]}{2+2\lambda_{x}\mathbb{E}[u_{1}(S_{x})]}$. Recall also that by Proposition 10.3, we have the following alternative expression for mean relevant work for the same arrival rate $\lambda_{x}$ and truncated job size $S_{x}$:
@@
Therefore, we can write the drift at speed 0 as $G\circ h_{k,x}(0,0)=\lambda_{x}\mathbb{E}[S_{x}^{2}]+kC_{k}(x,\lambda_{x})$. We are now ready to state the general version of the Rec-ISQ-$k$ lower bound. This is the second lower bound of Theorem 4.2.
See 10.6
@@
The proof of this theorem follows the same argument as the proof of Theorem 8.6. Let $\mathbb{E}_{\mathfrak{r}}[\cdot]$ denote the Palm expectation taken over the moments when the arbitrary arrival of jobs of size $x$ occurs according to the recycling stream.
Let $J_{x}(w,i):=h_{k,x}(w+x,\min(i+1/k,1))-h_{k,x}(w,i)$ denote the increase in the test function due to the arrival of a size-$x$ job. Then $\mathbb{E}_{\mathfrak{r}}[J_{x}(W_{x},I)]$ denotes the mean size of the recycling jump in stationarity. Thus, by Lemma 7.6, $\mathbb{E}[G\circ h_{k,x}(W_{x},I)]=\mathbb{E}[\text{Poisson drift}]+(\lambda-\lambda_{x})\mathbb{E}_{\mathfrak{r}}[J_{x}(W_{x},I)]$. By Lemma B.7 we can write the expected Poisson drift as,
$\mathbb{E}[\text{Poisson drift}]$ $=\lambda_{x}\mathbb{E}[S_{x}^{2}]+2\mathbb{E}[W_{x}](-1+\rho_{x})+kC_{k}(x,\lambda)\sum_{i=0}^{k}\frac{k-i}{k}\mathbb{P}(I^{r}=i/k)$
@@
$\mathbb{E}[W_{x}^{\text{arb-}ISQ}]$ $=\frac{\lambda_{x}E[S_{x}^{2}]}{2(1-\rho_{x})}+\dfrac{kC_{k}(x,\lambda_{x})(1-\rho_{\mkern 1.5mu\overline{\mkern-1.5mux\mkern-1.5mu}\mkern 1.5mu})}{2(1-\rho_{x})}+\frac{(\lambda-\lambda_{x})\mathbb{E}_{\mathfrak{r}}[J_{x}(W_{x},I)]}{2(1-\rho_{x})}$
$=\dfrac{\lambda_{x}\mathbb{E}[S_{x}^{2}]}{2(1-\lambda_{x}\mathbb{E}[S_{x}])}+\dfrac{\lambda_{x}\mathbb{E}[v_{1}(S_{x})]}{2+2\lambda_{x}\mathbb{E}[u_{1}(S_{x})]}\cdot\dfrac{1-\rho_{\bar{x}}}{1-\rho_{x}}+\frac{(\lambda-\lambda_{x})\mathbb{E}_{\mathfrak{r}}[J_{x}(W_{x},I)]}{2(1-\rho_{x})},$
where we use the definition of $C_{k}(x,\lambda)$ in the second equality. Recall the definition of an increasing speed queue. The current speed $i$ bounds the number of jobs in the system, which in turn bounds the total work, which consists of at most $ki$ jobs, each with size at most $x$. Therefore, for $0<i<1$, the system can only visit states $(w,i)$ where $w\leq kix$.